OpenStudy (anonymous):

What is the equation of the line that passes through (12, 4) and is perpendicular to the graph of y = –3/4 x – 2?

OpenStudy (anonymous):

Line perpendiular to the given line will have a slope of 4/3 therefore y = 4/3 x + b plug x=12 and y =4, to get b 4 = 4/3 * 12 + b b = -12 therefore answer is y = 4/3 x - 12